OpenStudy (anonymous):

What is the slope of the line whose equation is y -2/3 = 0? slope = -2/3 slope = 0 slope = 2/3

OpenStudy (anonymous):

add 2/3 to each side of the equation

OpenStudy (anonymous):

generel notation y=mx+n and m is the slope.

OpenStudy (anonymous):

slope = -2/3 slope = 0 slope = 2/3

OpenStudy (anonymous):

answer is 0 since the coefficient of x is 0